# LocoVR: Multiuser Indoor Locomotion Dataset in Virtual Reality

### v1

@@ Appendix I LocoVR dataset statistics
In this section, we describe the statistics of the trajectory data contained in LocoVR. We collected trajectory data from 31 participants in total, resulting in 7071 trajectory sequences after data preprocessing. Since we collected trajectories from two participants simultaneously, each participant’s trajectory was counted separately. We removed short trajectories (less than 2m or 2s) and poor motion tracking data in the data preprocessing phase.
@@
Figure 11 shows the distance distributions of the trajectories. The figure shows that more than half of the trajectories are longer than 4m. Since virtual rooms are usually smaller than 7m by 7m, the distribution is reasonable to assume daily movements in a room.
Figure 11 shows the travel time distribution of the trajectories. It shows that more than half of the trajectories are longer than 5s, which would be enough to learn locomotion in a single room.
Figure 12 shows the minimum distance between two participants in each trajectory. As can be seen, about 25% of the trajectories are within 1m of the other participant, and more than 70% are within 2m (See Figure 13). It indicates that many of the trajectories could be influenced by the trajectories of the other participants when they are in close proximity, as people typically consider how their behaviors might affect others when they are located close to other people. In the rest of the cases where the participants were at least 2m away from each other, there could still be social motion behaviors that involve passing through each other at a distance to respect other people’s personal space or taking a less direct route to the goal to avoid the risk of physical conflict with the other.
Figure 9: Number of trajectories collected in a scene. The percentile represents the cumulative percentage of the number of trajectories above a certain value.
Figure 10: Travel distance of trajectories. The percentile represents the cumulative percentage of the travel distance above a certain value.
Figure 11: Travel time of trajectories. The percentile represents the cumulative percentage of travel time above a certain value.
Figure 12: Minimum distance between two participants in each trajectory. The percentile represents the cumulative percentile of the minimum distance below a certain value.
@@
Appendix J Gap between VR and the Real
We observed gaps between people’s behaviors in the real scenes vs. the VR scenes. We speculate that the differences in behaviors could be attributed to the lower fidelity of SMPL avatars, which lack detailed facial expressions that can make interaction different from real-world interactions. The lack of physicality of the objects can also impact user behavior as people may walk through furniture, which is impossible in the real world. However, this is less frequent as prior VR research has shown Simeone et al. (2017). Also, these gaps can be addressed by adding user interaction cues, such as adding colliders to objects or face trackers, and we expect that future advances in VR-UI would help bridge the gap between VR and the real world.
In addition, there is a visual difference between scan-reconstructed images observed in HM3D Ramakrishnan et al. (2021) and the images captured in the real world by cameras. There would be difficulty in using the scan-reconstructed images for other real-world applications, such as the vision navigation of robots, due to the domain gap. However, we expect that the scan-reconstructed image could be utilized for pre-training the model before fine-tuning on the real data or utilized as the real images through domain adaptation James et al. (2019).

### v2

@@ Appendix I LocoVR dataset statistics
In this section, we describe the statistics of the trajectory data contained in LocoVR. We collected trajectory data from 32 participants in total, resulting in 7071 trajectory sequences after data preprocessing. Since we collected trajectories from two participants simultaneously, each participant’s trajectory was counted separately. We removed short trajectories (less than 2m or 2s) and poor motion tracking data in the data preprocessing phase.
I.1 Statistics on the data
@@
Figure 11 Path efficiency is a metric used to evaluate the complexity of a trajectory, defined as the ratio of the straight-line distance to the goal to the actual traveled distance. The mean path efficiency in LocoVR is 0.81, comparable to that reported in THOR-MAGNI (Schreiter et al., 2024). Notably, the data in THOR-MAGNI were collected in a single, controlled experimental setting designed to emphasize the influence of surrounding objects or other pedestrians. In contrast, our dataset was collected in diverse, naturally occurring home environments that were not controlled by experimenters.
Figure 13 shows the distance distributions of the trajectories. The figure shows that more than half of the trajectories are longer than 4m. Since virtual rooms are usually smaller than 7m by 7m, the distribution is reasonable to assume daily movements in a room. Figure 13 shows the travel time distribution of the trajectories. It shows that more than half of the trajectories are longer than 5s, which would be enough to learn locomotion in a single room.
Figure 15 shows the speed distribution of the trajectories. The mean value is around 0.8, indicating the walking speed is relatively low compared to the open outdoor space since the scale of home environments is smaller.
Figure 15 depicts the number of peaks observed in the speed history of individual trajectories. It shows that more than half of the trajectories exhibit two or more peaks, indicating that participants frequently change their speed on their way to the goal. This behavior is likely influenced by interactions with other participants or the narrow and complex indoor environment.
Figure 17 shows the minimum distance between two participants in each trajectory. It is shown that approximately 25% of the trajectories are within 1m of the other participant, and more than 70% are within 2m (See Figure 18). It indicates that many of the trajectories could be influenced by the trajectories of the other participants when they are in close proximity, as people typically consider how their behaviors might affect others when they are located close to other people. In the rest of the cases where the participants were at least 2m away from each other, there could still be social motion behaviors that involve passing through each other at a distance to respect other people’s personal space or taking a less direct route to the goal to avoid the risk of physical conflict with the other.
Figure 17 illustrates the distribution of closing speeds between participants. The relatively high closing speeds, considering the room scale, suggest that participants need to remain attentive to the movements of others.
Figure 10: Number of trajectories in each scene
Figure 11: Path efficiency (straight-line distance to the goal/traveled distance) in each trajectory.
Figure 12: Travel distance in each trajectory
Figure 13: Travel time in each trajectory
Figure 14: Speed of participants.
Figure 15: Number of peaks in speed history in each trajectory
Figure 16: Minimum distance between participants in each trajectory.
Figure 17: Closing speed between participants.
@@
I.2 Statistics on the participants
There were 32 participants in total, comprising 21 males and 11 females, with ages ranging from 18 to 42. From this pool, pairs were formed to conduct 25 experiments, each involving a unique pair (Table12). The experiments included various combinations of male-male, female-female, and male-female pairs, as well as pairs of friends and nonfriends, as shown in the Table13. As reactions between pairs in close proximity could influenced by attributes and interpersonal relationships, further data analysis may provide new insights into the relationship between these attributes, relationships, and behavioral patterns. It could be an intriguing study from the perspective of cognitive and behavioral sciences.
Table 12: Participants demographics
$Age$ $Number\;of\;participants$
$Male$ $Female$
Under 20 5 3
20 to 29 15 8
Over 30 1 0
Table 13: Diversity of pairs
$Relationship$ $Number\;of\;participants$
$Male-Male$ $Female-Female$ $Male-Female$
$Friends$ 2 2 5
$Nonfriends$ 9 1 6
I.3 Why two person ? - Motivation of two-person experiment
In contrast to conventional studies that focus on crowd dynamics in open public spaces, our research emphasizes room-scale social motion behaviors between two individuals, particularly in confined indoor spaces. For example, interpersonal behaviors such as proxemics, trajectory negotiation, and mutual space adaptation are affected and induced by the narrow geometry. We consider this is a fundamental study on the room-scale social motion behaviors, and believe this focus does not limit the dataset’s utility. Researchers can build upon it to study two-person interactions in isolation or as a basis for modeling interactions in more complex, multi-person environments. In addition, more than two-person interaction could be considered as a rare scenario in home environments. For instance, recent census data indicates that in the US, 60% of households consist of two people or fewer, and 80% consist of three people or fewer. Also, indoor home spaces are generally smaller than open public spaces, and activities in such settings are often more individualized. As a result, residents typically move independently rather than engaging in collaborative or coordinated movements within the same space. This relevance underscores the practical utility of our dataset for studying interaction dynamics that are directly applicable to real-world scenarios.
Appendix J Influence of gap between VR and the Real
We consider the virtual avatar an effective tool for addressing the VR/real-world gap. Additionally, we introduced a filter to exclude data exhibiting inappropriate behaviors. In the following sections, we discuss the impact of this gap.
(1) In our experiment, participants were aware that the virtual avatars were synchronized with real humans sharing the same physical space. Also, the avatar enables participants to percept relative position between their body and surrounding objects. These awareness discouraged socially or physically inappropriate behavior, mitigating the potential impact of the VR/real gap, as demonstrated in recent studies on VR locomotionYun et al. (2024); Simeone et al. (2017). In addition, we have introduced a filter to detect instances of users passing through virtual objects to remove such data from the dataset.
(2) Our evaluation used locomotion data collected in physical spaces as test data. Models trained on the LocoVR dataset outperformed those trained on other physically collected datasets (GIMO/THOR-MAGNI), demonstrating that VR-collected data is effective when applied to real-world scenarios.
For the futurework, more expressive avatar to facilitate non-verbal communication during walk, or user interaction cues such as haptic devices could further reduce the gap between VR/Real. We anticipate that future advancements in VR technologies will further contribute to bridging this gap.